OpenStudy (anonymous):

how do you write log5 625=y in exponential form?

OpenStudy (anonymous):

\[\log _{5}625=y\]

OpenStudy (anonymous):

\[\log _{b}y=x\]becomes\[b^{x}=y\]

OpenStudy (anonymous):

5^y = 625

OpenStudy (anonymous):

I would first rewrite 625 as a power of 5.

OpenStudy (anonymous):

Oh, you're not solving it. Nvm.